13.

A total of 19,000 students entered Big State University in the fall of 2013. The following table breaks down the student body by race and gender: White African-American Hispanic Asian Other Total Male Female Total 2,8503,8006,650 1,520 3,420 1,9002,850 1,8242,736 4,560 9,234 9,766 19,000 570 3,040 950 950 380 Complete the following table by choosing the correct value representing the portion of students that is both female and Asian in fraction, decimai, and percent form Fraction Decimal Percent Female and AsianComplete the following table by choosing the correct value representing the portion of students that is both female and Asian in fraction, decimai, and percent form. Fraction Decimal Percent Female and Asian The portion of students that are female and Hispanic is 9300 Which of the following represent fractions that are equivalent to? Check all that apply 38 19,000 38 1.900 76 3,800 380Complete the following table by choosing the correct value representing the portion of students that is both female and Asian in fraction, decimal, and percent form Fraction Decimal Percent Female and Asian 1/40 1/20 1/10 1/5 380 female and Hispanic is R8? Check all that The portion of student apply 19.000 Which of the following represent fractions that are equivalent to 1900 19,000 38 19,000 38 1.900 76 3,800 380Complete the following table by choosing the correct value representing the portion of students that is both female and Asian in fraction, decimai, and percent form. Fraction Decimal Percent Female and Asian 0.025 banic is19,00 anic is 13800. Which of the following represent fractions that are equivalent to? Check all that The portion of students that are female apply 0.05 0.1 38 19,000 0.2 38 1.900 76 380Complete the following table by choosing the correct value representing the portion of students that is both female and Asian in fraction, decimal, and percent form. Fraction Decimal Percent Female and Asian 2.5% 380 The portion of students that are female and Hispanic is hich of the following represent fractions that are equivalent to 19000 Check all that apply 590 38 19,000 38 76 3,800 380
Answer #1
Fraction Decimal Percent
Female and Asian (1900/19000)=1/10 0.1 (0.1)(100)= 10%

The portion of students that are female and Hispanic is 380/19000.

The fractions 38/1900 and 76/3800 are equivalent to 380/19000 because they are equal to 0.02.

Hence, option(B) and (C) are correct.